# SeFi-IDE: Semantic-Fidelity Identity Embedding for Personalized Diffusion-Based Generation

### 31 Jan 2024 v1

@@ Abstract
00footnotetext: $*{}^{*}$ indicates equal contribution.00footnotetext: indicates corresponding author.
@@
Advanced diffusion-based Text-to-Image (T2I) models, such as the Stable Diffusion Model, have made significant progress in generating diverse and high-quality images using text prompts alone. However, T2I models are unable to accurately map identities (IDs) when non-famous users require personalized image generation. The main problem is that existing T2I models do not learn the ID-image alignments of new users. The previous methods either failed to accurately fit the face region or lost the interactive generative ability with other existing concepts in T2I models (i.e., unable to generate other concepts described in given prompts such as scenes, actions, and facial attributes). In this paper, we focus on accurate and semantic-fidelity ID embedding into the Stable Diffusion Model for personalized generation. We address this challenge from two perspectives: face-wise region fitting, and semantic-fidelity token optimization. Specifically, we first visualize the attention overfit problem, and propose a face-wise attention loss to fit the face region instead of the whole target image. This key trick significantly enhances the ID accuracy and interactive generative ability with other existing concepts. Then, we optimize one ID representation as multiple per-stage tokens where each token contains two disentangled features. This expansion of the textual conditioning space enhances semantic-fidelity control. Extensive experiments validate that our results exhibit superior ID accuracy and manipulation ability compared to previous methods. More results: https://com-vis.github.io/SeFi-IDE/.
@@
Figure 1: The face identity (ID) embedding of previous personalized generation methods has two problems:(1) Attention Overfit : The attention of Textural Inversion [8] and ProSpect [48] towards to the “$*V*$” takes over the whole image, disrupting the interaction between newly embedded concepts and other existing concepts. Consequently, their attention towards to the “cup” is wrong, which results in the failure of the given prompt. (2) Limited Semantic-Fidelity: Despite alleviating overfit, Celeb Basis [47] introduces excessive face prior, limiting the semantic-fidelity of the learned ID embedding. Specifically, this means that the “cup” attention map continues to concentrate on the face region. We propose Face-Wise Attention Loss (Sec. 3.1) and Semantic-Fidelity Token Optimization (Sec. 3.2) to address problem (1) and (2) respectively.
1 Introduction
@@
Recently, Text-2-Image (T2I) models, such as the Stable Diffusion Model [29], have demonstrated an impressive ability to generate diverse, high-quality, and semantic-fidelity images using text prompts alone, thanks to image-aligned language encoders [25] and diffusion-based generative models [5, 21]. However, the challenge of personalized generation still remains, because the accurate person-specific face manifold can not be represented by text tokens, especially for the non-famous users whose data are not included in the training dataset. In this paper, we focus on accurate and semantic-fidelity identity embedding for personalized diffusion-based generation using one face image.
@@
The previous methods have two problems require to be addressed: (1) Their fine-tuning strategies [8, 19, 30] tend to fit the whole target image rather than the face region. This resulted in the low ID accuracy and the difficulty of generating other concepts in the given prompt, such as ID-unrelated scenes (e.g., Eiffel Tower), ID-related facial attributes (e.g., expressions and age), and actions (e.g., playing the guitar). Particularly for actions, it is more challenging to generate prompt-fidelity human motions and human-object interactions, which can be shown in Fig. 1. (2) Their ID embedding methods lack the semantic-fidelity representations for facial attributes, which results in that human faces are treated as objects without non-rigid and diverse deformations. Although the method such as Celeb Basis [47] can achieve an accurate ID mapping, it is unable to manipulate the facial attributes of the target image, such as expressions.
@@
To address these problems, we propose our identity embedding method from two perspectives: (1) Face-Wise Region Fit: We first visualize the attention overfit problem of the previous methods from the attention feature activation maps, and then propose a face-wise attention loss to fit the face region instead of the whole target image. This key trick can improve the ID accuracy and interactive generative ability with the existing concepts in the original Stable Diffusion Model. (2) Semantic-Fidelity Token Optimization: We optimize one ID representation as several per-stage tokens and each token consists of two disentangled features. This approach expands the textual conditioning space and allows for semantic-fidelity control ability. Our extensive experiments confirm that our method achieves higher accuracy in ID embedding and produces a wider range of scenes, facial attributes, and actions compared to previous methods.
@@
  • Moreover, our embedding method has the potential to be applied to other categories, including dogs, due to its independence from any prior facial knowledge.
@@
Figure 2: The overview of our framework. We first propose a novel Face-Wise Attention Loss (Sec. 3.1) to alleviate the attention overfit problem and make the ID embedding focus on the face region to improve the ID accuracy and interactive generative ability. Then, we optimize the target ID embedding as five per-stage tokens with disentangled features to expend textural conditioning space with semantic-fidelity control ability (Sec. 3.2).
2 Related Work
2.1 Text-Based Image Synthesis and Manipulation
Synthesis. Previous models such as GAN [16, 10, 45, 44, 11, 43], VAE [18, 37, 46, 42], Autoregressive [27, 7], Flow [26, 6] were adopted to model the dataset distribution. Recent advanced diffusion models [5, 21] have shown excellent diversity and fidelity in text-to-image synthesis [22, 14, 27, 29, 31, 2]. Conditioned on the text embedding of the text encoder [25], these diffusion-based models are optimized by a simple denoising loss and can generate a new image by sampling Gaussian noise and a text prompt.
Manipulation. However, these diffusion-based generative models have difficulties in generating novel concepts [3] or controlling the identity of generated objects [8]. Existing methods either directly blended the latent code of objects [1] to the generated background, or failed to understand the scenes correctly [41], which results in the obvious artifacts. To further solve this problem, some works [12, 20, 23, 24] adopted attention-based methods to manipulate target objects, but cannot balance the trade-off between content diversity and identity accuracy.
2.2 Personalized Generation of T2I Models
@@
The existing diffusion-based T2I models are limited by generating ID-accurate images for the specific unseen concepts. Thus, using images from the novel concepts for fine-tuning enables us to obtain a personalized model, which can insert novel concepts into the original model and synthesize ID-specific new scenes, appearances, and actions. Inspired by the GAN Inversion [39], recent diffusion-based personalized generation works can be divided into two categories: (1) Fine-Tuning T2I model: DreamBooth [30] fine-tunes all weight of the T2I model on a set of images with the same ID and marks it as the specific token. (2) Token Optimization: The works such as Textural Inversion [8], ProSpect [48], and Celeb Basis [47] optimize the text embedding of special tokens to map the specific ID into the T2I model, where the T2I model is fixed in the optimization process. Recently, these two strategies are combined for manipulating multi-concept interactions [19] or saving fine-tuning time and parameter amount [33, 35].
@@
ID Embedding for Faces. Previous methods [32, 9] try to train a inversion encoder for face embedding, but face ID-oriented mapping is difficult to be obtained from a naively optimized encoder. Moreover, fine-tuning the T2I model on large-scale images often causes concept forgetting. Towards this, Celeb Basis [47] adopt a pre-trained face recognition model and a face ID basis to obtain an ID representation for one single face image and Face0 [36] learned to project the embeddings of recognition models to context space of Stable Diffusion. However, introducing too strong face prior makes it difficult to manipulate diverse facial attributes and fails to generalize to other concept embedding.
3 Method
@@
Embedding one specific ID into the Stable Diffusion Model for personalized generation using only one single face image has two technical requirements: accuracy and semantic-fidelity. As shown in Fig. 2, we propose our ID embedding pipeline from two key perspectives: (1) Face-Wise Attention Loss: Towards the improvements of ID accuracy and interactive generative ability with the existing concepts in the original model, we propose a face-wise attention loss in Sec. 3.1. (2) Semantic-Fidelity Token Optimization: For diverse manipulation, we optimize one ID representation as several per-stage tokens, and each token consists of two disentangled features, which can be seen in Sec. 3.2.
3.1 Face-Wise Attention Loss
@@
We first analyze and visualize the attention fit problem of previous methods. Then, we present an accessible prior from the Stable Diffusion Model to improve the interactive generative ability.
@@
Image Fit vs. Face-Region Fit. Previous methods rely on learning from multiple images of a target object to grasp relevant features, such as Textual Inversion [8] and DreamBooth [30]. However, when only a single image is available, they are prone to fitting to the whole target image and the learned embedding tokens tend to influence regions beyond the face region during the cross-attention stages. As a result, they lack the good interactive generative ability with the existing concepts in the original model [35]. In other words, during the inference, the generated results from the personalized model may not be consistent to the text prompts. For example, as shown in Fig. 1, the given prompt is “a V* is enjoying a cup of latte”, but the methods with attention overfit problem fail to generate the “cup” content. The same problem can also be seen in Fig. 3, which given some facial attribute such as “old” in the prompt, this diffusion-based generation process just fails. Our ID embedding optimization can focus on the ID-related face regions and neglect the unrelated background features, which can improve the ID accuracy and interactive generative ability.
@@
Make Best of Stable Diffusion Model. Multiple target images are necessary for previous methods to acquire concept-related embedding tokens. These images enable users to utilize text prompts to manipulate different object poses, appearances, and contextual variations. One target image fails to achieve such generalization. However, Stable Diffusion Model has learnt a strong general concept prior for various sub-categories. For example, different human identities belong to the general concept “person”, and different dog categories such as Corgi and Chihuahua belong to the general concept “dog”. Therefore, it is reasonable to adopt this prior knowledge to achieve one-shot learning. To meet our more higher requirements, we aim to have flexibility to manipulate the ID-specific regions of final images. In other words, when we want to generate images corresponding to “a photo of V is playing guitar”, only handling portrait or face image generation is not enough for this prompt. Therefore, we adopt “the face of person” as our general concept prior for ID embedding, because when provided with the prompt “a photo of the face of person”, Stable Diffusion Model can generate an image of a person with a randomly assigned identity and constrain the region where the generated person appears in the final image.
@@
We propose to use a reference prompt $P_{r}$ that remains consistent with the general concept of different IDs, which replaces the placeholder word (“V”) with “person” in “a photo of the face of V”. Then, we use this attention map derived from $P_{r}$ as a constraint to restrict the attention corresponding to the placeholder word (“V*”) in the target prompt “a photo of the face of V*”. This approach allows the ID embedding to focus on the face region associated with the target ID while maintaining the coherence of the general concept. Specifically, we first utilize CLIP encoder [31] to encode the reference prompt and target prompt as $\bm{P_{r}}$ and $\bm{P_{t}}$. The $\bm{P_{t}}$ is fed to the self-attention module to obtain five per-stage tokens $\bm{P}={\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. Then, we adopt text encoder [31] to obtain their corresponding key ($\bm{K}$) and value ($\bm{V}$) features ${\bm{K_{r}},\bm{V_{r}}}$ and ${\bm{K_{t}},\bm{V_{t}}}$. Then, each of the K features are send to the cross-attention module to calculate the attention map $\bm{A_{r}}$ and $\bm{A_{t}}$ with latent image features $\bm{Q}$ respectively. The $\bm{A_{t}}$ is constrained by the $\bm{A_{r}}$ within the corresponding concept representation as follows:
$\mathcal{L}_{attention}=||\bm{A_{r}}-\bm{A_{t}}||_{2}^{2}$ (1)
3.2 Semantic-Fidelity Token Optimization
We first present the disadvantages of previous methods. And then we introduce our optimization strategy.

### 22 Mar 2024 v2

# Beyond Inserting: Learning Identity Embedding for Semantic-Fidelity Personalized Diffusion Generation

@@ Abstract
00footnotetext: $*{}^{*}$ indicates equal contribution and indicates corresponding author. Yang Li, Songlin Yang, Wei Wang, and Jing Dong are with the Center for Research on Intelligent Perception and Computing (CRIPAC), State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS), Institute of Automation, Chinese Academy of Sciences (CASIA), Beijing, 100190, China. Yang Li and Songlin Yang are also with the School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, 100049, China. (E-mail: liyang2022@ia.ac.cn; yangsonglin2021@ia.ac.cn; wwang@nlpr.ia.ac.cn; jdong@nlpr.ia.ac.cn)
@@
Advanced diffusion-based Text-to-Image (T2I) models, such as the Stable Diffusion Model, have made significant progress in generating diverse and high-quality images using text prompts alone. However, when non-famous users require personalized image generation for their identities (IDs), the T2I models fail to accurately generate their ID-related images. The main problem is that pre-trained T2I models do not learn the mapping between the new ID prompts and their corresponding visual content. The previous methods either failed to accurately fit the face region or lost the interactive generative ability with other existing concepts in T2I models. In other words, they are unable to generate T2I-aligned and semantic-fidelity images for the given prompts with other concepts such as scenes (“Eiffel Tower”), actions (“holding a basketball”), and facial attributes (“eyes closed”). In this paper, we focus on inserting accurate and interactive ID embedding into the Stable Diffusion Model for semantic-fidelity personalized generation. We address this challenge from two perspectives: face-wise region fitting and semantic-fidelity token optimization. Specifically, we first visualize the attention overfit problem and propose a face-wise attention loss to fit the face region instead of entangling ID-unrelated information, such as face layout and background. This key trick significantly enhances the ID accuracy and interactive generative ability with other existing concepts. Then, we optimize one ID representation as multiple per-stage tokens where each token contains two disentangled features. This expansion of the textual conditioning space improves semantic-fidelity control. Extensive experiments validate that our results exhibit superior ID accuracy, text-based manipulation ability, and generalization compared to previous methods.
Index Terms:
Generative Models, Text-to-Image Generation, Diffusion Models, and Personalized Generation
@@
Figure 1: Previous methods for inserting new identities (IDs) into pre-trained Text-to-Image diffusion models for personalized generation have two problems: (1) Attention Overfit : As shown in the activation maps of Textural Inversion [1] and ProSpect [2], their “V*” attention nearly takes over the whole images, which means the learned embeddings try to encode both the human faces and ID-unrelated information in the reference images, such as the face region layout and background. This problem extremely limits their generative ability and disrupts their interaction with other existing concepts such as “cup”, which results in the failure of the given prompt (i.e., they fail to generate the image content aligned with the given prompt). (2) Limited Semantic-Fidelity: Despite alleviating overfit, Celeb Basis [3] introduces excessive face prior, limiting the semantic-fidelity of the learned ID embedding (e.g., the “cup” attention still continues to the “V*” face region and this limitation hinders the control of facial attributes such as “eyes closed”). Therefore, we propose Face-Wise Region Fit (Sec. III-B) and Semantic-Fidelity Token Optimization (Sec. III-C) to address problem (1) and (2) respectively. More results: https://com-vis.github.io/SeFi-IDE/.
I Introduction
@@
Recently, Text-to-Image (T2I) models, such as the Stable Diffusion Model [4], have demonstrated an impressive ability to generate diverse, high-quality, and semantic-fidelity images using text prompts alone, thanks to image-aligned language encoders [5] and diffusion-based generative models [6, 7]. However, the challenge of personalized generation still remains, because the accurate person-specific face manifold can not be represented by text tokens, especially for the non-famous users whose data are not included in the training dataset. In this paper, we focus on learning the accurate identity embedding for semantic-fidelity personalized diffusion-based generation using only one face image.
@@
The previous methods for this task have two problems that need to be addressed: (1) Attention Overfit: Their fine-tuning strategies [8, 9], such as Texural Inversion [1] and ProSpect [2], tend to fit the whole target image rather than the ID-related face region, which entangle face layout and background information into the ID embedding. This results in the low ID accuracy and the difficulty of generating other existing concepts in the given prompt, such as ID-unrelated scenes (e.g., “Eiffel Tower”), ID-related facial attributes (e.g., expressions and age), and actions (e.g., “holding a basketball”). Particularly for actions, it is more challenging to generate prompt-fidelity human motions and human-object interactions, which can be shown in Fig. 1. (2) Limited Semantic-Fidelity: Their ID embedding methods lack the semantic-fidelity representations for facial attributes, which results in that human faces are treated as objects without non-rigid and diverse deformations. Although Celeb Basis [3] can achieve an accurate ID mapping, it is unable to manipulate the facial attributes of the target image, such as expressions (e.g., “eyes closed” in Fig. 1).
@@
To address these problems, we propose our identity embedding method from two perspectives: (1) Face-Wise Region Fit: We first visualize the attention overfit problem of the previous methods from the attention feature activation maps and then propose a face-wise attention loss to fit the face region instead of the whole target image. This key trick can improve the ID accuracy and interactive generative ability with the existing concepts in the original Stable Diffusion Model. (2) Semantic-Fidelity Token Optimization: We optimize one ID representation as several per-stage tokens, and each token consists of two disentangled features. This approach expands the textual conditioning space and allows for semantic-fidelity control ability. Our extensive experiments validate that our method achieves higher accuracy in ID embedding and is able to produce a wider range of scenes, facial attributes, and actions compared to previous methods.
@@
  • Our method does not rely on any prior facial knowledge, which has the potential to be applied to other categories.
@@
Figure 2: The overview of our framework. We first propose a novel Face-Wise Attention Loss (Sec. III-B) to alleviate the attention overfit problem and make the ID embedding focus on the face region to improve ID accuracy and interactive generative ability. Then, we optimize the target ID embedding as five per-stage tokens pairs with disentangled features to expend textural conditioning space with semantic-fidelity control ability (Sec. III-C).
II Related Work
II-A Text-Based Image Synthesis and Manipulation
Previous models such as GAN [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], VAE [20, 21, 22, 23], Autoregressive [24, 25], Flow [26, 27] were adopted to model the dataset distribution, and then synthesize new realistic images through sampling from the modeled distribution. Based on these, text-driven image manipulation [28, 29, 30] has achieved significant progress using GANs by combining text representations such as CLIP [5]. These methods work well on structured scenarios (e.g. human face editing), but their performance in fine-grained multi-modal alignment is not very satisfactory. Recent advanced diffusion models [6, 7] have shown excellent diversity and fidelity in text-to-image synthesis [31, 32, 24, 4, 33, 34]. Conditioned on the text embedding of the text encoder [5], these diffusion-based models are optimized by a simple denoising loss and can generate a new image by sampling Gaussian noise and a text prompt. Thanks to the powerful capabilities of diffusion in T2I generation, works [35, 36, 37] achieve state-of-the-art text based image editing quality over diverse datasets, often surpassing GANs. Although most of these approaches enable various global or local editing of an input image, all of them have difficulties in generating novel concepts [38] or controlling the identity of generated objects [1]. Existing methods either directly blended the latent code of objects [39, 40] to the generated background, or failed to understand the scenes correctly [41], which results in the obvious artifacts. To further solve this problem, some work [42, 43, 44, 45] adopted attention-based methods to manipulate target objects, but fail to balance the trade-off between content diversity and identity accuracy.
II-B Personalized Generation of Diffusion-Based T2I Models
@@
Using images from the new concepts for fine-tuning can obtain a personalized model, which can insert new concepts into the original model and synthesize concept-specific new scenes, appearances, and actions. Inspired by the GAN Inversion [46], recent diffusion-based personalized generation works can be divided into three categories: (1) Fine-Tuning T2I model: DreamBooth [9] fine-tunes all weight of the T2I model on a set of images with the same ID and marks it as the specific token. (2) Token Optimization: Textual Inversion [1], ProSpect [2], and Celeb Basis [3] optimize the text embedding of special tokens to map the specific ID into the T2I model, where the T2I model is fixed in the optimization process. (3) Tuning Free: ELITE [47] learning an encoder to customize a visual concept provided by the user without further fine-tuning. BootPIG [48] follows a bootstraping strategy by utilizing a pre-trained U-Net model to steer the personalization generation. Except for those, Token Optimization and Fine-Tuning are combined to manipulate multi-concept interactions [8] or saving fine-tuning time and parameter amount [49, 50, 51, 52].
@@
ID Embedding for Faces. Previous methods [53, 54] try to train an inversion encoder for face embedding, but face ID-oriented mapping is difficult to be obtained from a naively optimized encoder. Moreover, fine-tuning the T2I model on large-scale images often causes concept forgetting. For this, Celeb Basis [3] adopts a pre-trained face recognition model and a face ID basis to obtain an ID representation for one single face image, and Face0 [55] learned to project the embeddings of recognition models to the context space of Stable Diffusion. Except for ID representation, FaceStudio [56] deployed a CLIP vision encoder [5] to extract the structure features. InstantID [57] handled image generation in various styles by designing a learnable IdentityNet to grasp strong semantics. However, introducing too strong face prior makes it difficult to manipulate diverse facial attributes and fails to generalize to other concept embedding. FastComposer [58] used a delayed subject conditioning strategy to avoid subject overfitting, but they only focus on faces and fail to interact with other objects such as “sofa” as shown in Fig. 9. While PhotoMaker [59] proposed an ID-oriented dataset that includes diverse scenarios and fine-tuning part of the Transformer [60] layers in the image encoder to mitigate contextual information loss. Nevertheless, the training of Transfromer will sacrifice the compatibility with existing pretrained community models.
III Method
@@
Embedding one new identity (ID) into the Stable Diffusion Model for personalized generation using only one single face image has three technical requirements: accuracy, interactivity, and semantic-fidelity. Our learned ID embedding focuses on the face region and adopts disentangled token representation, which has flexible face spatial layout, interactive generation ability with existing concepts (e.g., generating interaction motion with other objects), and fine-grained manipulation ability (e.g., editing the facial expressions). This means that our method improves both ID accuracy and manipulation ability. As shown in Fig. 2, we propose our ID embedding pipeline from two key perspectives: (1) Face-Wise Attention Loss: Towards the improvements in ID accuracy and interactive generative ability with existing concepts in the original model, we propose a face-wise attention loss in Sec. III-B. (2) Semantic-Fidelity Token Optimization: For diverse manipulation, we optimize one ID representation as several per-stage tokens, and each token consists of two disentangled embeddings, which can be seen in Sec. III-C. In the following sections, we first give an introduction of the pre-trained Stable Diffusion Model [4], and we then provide the details of our method.
III-A Preliminary
Diffusion-Based T2I Generation. Our utilized Stable Diffusion Model [4] consists of a CLIP text encoder [5], an AutoVAE [61] and a latent U-Net [62] module. Given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ ($H$ and $W$ represent the size of target image), the VAE encoder $\mathcal{E}(\cdot)$ maps it into a lower dimensional latent space as $\bm{z}=\mathcal{E}(\bm{x})\in{\mathbb{R}^{h\times w\times c}}$ followed by a corresponding decoder $\mathcal{D}(\cdot)$ to map the latent vectors back as $\mathcal{D}(\mathcal{E}(\bm{x}))\approx\bm{x}$. The $h,w$ and $c$ are the dimensions of latent tensor $\bm{z}$. Given any user provided prompts $y$, the tokenizer of the CLIP text encoder $\mathcal{E}_{text}(\cdot)$ divides and encodes $y$ into integer tokens. Correspondingly, by looking up the dictionary, a word embedding group $\bm{Y}$ can be obtained. Then, the CLIP text transformers $\mathcal{T}_{text}$ encode $\bm{Y}$ to generate text condition vectors $\mathcal{T}_{text}(\bm{Y})$, which serve as a condition to guide the training of the latent U-Net denoiser $\epsilon_{\theta}(\cdot)$:
$\mathcal{L}=\mathbb{E}_{\bm{z}\sim\mathcal{E}(\bm{x}),\bm{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t,\mathcal{T}_{text}(\bm{Y}))\right\|\right],$ (1)
where $\bm{\epsilon}$ denotes for the unscaled noise and $t$ is the timestep. $\bm{z}_{t}$ is the latent variable of a forward Markov chain $\mathrm{s.t.}\bm{z}_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon}$, where $\alpha\in\left[0,1\right]$ is a hyper-parameter that modulates the quantity of noise added. Given a latent noise vector $\bm{z}_{t}$ in the timestep $t$, the model learns to denoise it to $\bm{z}_{t-1}$. During inference, a random Gaussian latent vector $\bm{z}_{T}$ is iteratively denoised to $\bm{z}_{0}$.
Figure 3: The details of text condition and K-V feature implementation differences.
Cross-Attention for Text Condition. As shown in the upper block of Fig. 3, the text prompt is first tokenized to a word embedding group $\bm{Y}$, and then encoded by the text transformers to generate text condition $\mathcal{T}_{text}(\bm{Y})$. Given the latent image features $\bm{f}$, the cross attention operation updates the latent features as:
$\bm{Q}=\bm{W^{q}}\bm{f},\bm{K}=\bm{W^{k}}\mathcal{T}_{text}(\bm{Y}),\bm{V}=\bm{W^{v}}\mathcal{T}_{text}(\bm{Y}),$ (2)
$Attention(\bm{Q},\bm{K},\bm{V})=Softmax(\frac{\bm{Q}\bm{K^{T}}}{\sqrt{d}})\bm{V}.$ (3)
where $\bm{W^{q}}$, $\bm{W^{k}}$, and $\bm{W^{v}}$ map the inputs to Query, Key, and Value features, respectively. The $d$ is the output dimension of Key and Query features.
Previous work has shown that the CLIP text embedding space is expressive enough to capture image semantics [1, 2]. Specifically, a placeholder string, “$V^{*}$”, is designated in the prompt $y$ to represent the identity-related feature we wish to learn. During the word embedding process, the vector associated with the word “$V^{*}$” is replaced by the learned ID embedding $\bm{P}$. Thus, we can combine “$V^{*}$” with other words to achieve personalized creation. In this work, we focus on learning accurate and interactive ID embedding $\bm{P}$.
Algorithm 1 Calculating the Face-Wise Attention Loss
0:  Query features $\mathbf{Q}$, reference key features $\bm{K_{r}}$, target
  key features $\bm{K_{t}}$, attention score modules ATT in cross-
  attention blocks, and attention map rearrange and resize
  function Re2.
  $\mathcal{L}_{attention}$
  for $i\leftarrow 1$ to $S$ do
     $\bm{A_{r}}.\text{append}(\mathbf{\textbf{ATT}_{i}}(\bm{Q},\bm{K_{r}}))$
     $\bm{A_{t}}.\text{append}(\mathbf{\textbf{ATT}_{i}}(\mathbf{Q},\bm{K_{t}}))$
  end for
  Rearrange and resize the attention map before calculating the loss
  $\bm{A_{r}}=\textbf{Re2}(\bm{A_{r}})\in\mathbb{R}^{8\times 77\times{32}\times{32}}$
  $\bm{A_{t}}=\textbf{Re2}(\bm{A_{t}})\in\mathbb{R}^{8\times 77\times{32}\times{32}}$
  Calculate the Face-Wise Attention Loss
  $\mathcal{L}_{attention}=\text{MSE}(\bm{A_{t}},\bm{A_{t}})$
  return  $\mathcal{L}_{attention}$
III-B Face-Wise Attention Loss
@@
We first analyze and visualize the attention overfit problem of previous methods. Then, we present an accessible prior from the Stable Diffusion Model, instead of the face prior from other models such as face recognition models, to improve both the embedding accuracy and interactive generative ability at the same time. These are our motivations to propose our Face-Wise Attention Loss. Finally, we present the details of the loss implementations.
Figure 4: The different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens. (1) Progressively Adding: We add different ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$ tokens to the conditioning information in ten steps. We found that the initial tokens effect more the layout of generation content (e.g., face region location, and poses), while the latter tokens effect more the ID-related details. (2) Progressively Substituting: We then substitute different $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens of ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. We found that $\bm{P_{i}^{V}}$ contribute to the vast majority of ID-related conditioning information, and the $\bm{P_{i}^{K}}$ contribute more to textural details, such as environment lighting.
@@
Image Fit vs. Face-Region Fit. Previous methods rely on learning from multiple images of a target object to grasp relevant features, such as Textual Inversion [1] and DreamBooth [9]. However, when only a single image is available, they are prone to fitting to the whole target image (including ID-unrelated face layout and background information), and the learned embedding tends to influence regions beyond the face region during the cross-attention stages. As a result, they lack the interactive generative ability with the existing concepts in the original model [50]. In other words, during the inference, the generated results from the personalized model may not be consistent to the text prompts. For example, as shown in Fig. 1, the given prompt is “a $V^{*}$ is enjoying a cup of latte”, but the methods with attention overfit problem fail to generate the “cup” content. The same problem can also be seen in Fig. 7, which given some facial attributes such as “old” in the prompt, the diffusion-based generation process just fails. Our ID embedding optimization can focus on ID-related face regions and neglect the unrelated background features, which can simultaneously improve the ID accuracy and interactive generative ability.
@@
Make Best of Stable Diffusion Model. Multiple target images are necessary for previous methods to acquire concept-related embedding. These images allow users to use text prompts to manipulate different poses, appearances, and contextual variations of objects. One target image fails to achieve this generalization. However, Stable Diffusion Model has learnt a strong general concept prior for various sub-categories. For example, different human identities belong to the general concept “person”, and different dog categories such as Corgi and Chihuahua belong to the general concept “dog”. Therefore, it is reasonable to adopt this prior knowledge to achieve one-shot learning. To meet our more higher requirements, we aim to have flexibility to manipulate the ID-specific regions of final images. In other words, when we want to generate images corresponding to “a photo of $V^{*}$ is playing guitar”, only handling portrait or face image generation is not enough for this prompt. Therefore, we adopt “the face of person” as our general concept prior for ID embedding, because when provided with the prompt “a photo of the face of person”, Stable Diffusion Model can generate a face image of a person with a randomly assigned identity and constrain the region where the generated person appears in the final image.
@@
Specifically, we propose to use a reference prompt $y_{r}$ that remains consistent with the general concept of different IDs, which replaces the placeholder word (“$V^{*}$”) with “person” in prompts (i.e., $y_{r}:=$“a photo of the face of person”). Then, we use this attention map derived from $y_{r}$ as a constraint to restrict the attention corresponding to the placeholder word (“$V^{*}$”) in the target prompt “a photo of the face of $V^{*}$”. This approach allows the ID embedding to focus on the face region associated with the target ID while maintaining the coherence of the general concept. Specifically, we first embed the reference prompt and target prompt as word embedding groups $\bm{Y_{r}}$ and $\bm{Y_{t}}$. The ID embedding $\bm{P}$ in $\bm{Y_{t}}$ is fed into a self-attention module to obtain per-stage token embeddings ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$, which will be introduced in the subsequent section. Then, we adopt text encoder transformers $\mathcal{T}_{text}$ to obtain their corresponding key ($\bm{K}$) and value ($\bm{V}$) features ${\bm{K_{r}},\bm{V_{r}}}$ and ${\bm{K_{t}},\bm{V_{t}}}$. Then, each of the K features are send to the cross-attention module to calculate the attention map $\bm{A_{r}}$ and $\bm{A_{t}}$ with latent image features $\bm{Q}$ respectively. The $\bm{A_{t}}$ is constrained by the $\bm{A_{r}}$ within the corresponding representation of the concept as follows:
$\mathcal{L}_{attention}=||\bm{A_{r}}-\bm{A_{t}}||_{2}^{2}.$ (4)
The detailed Face-Wise Attention Loss computation pipeline is depicted in Algorithm 1.
III-C Semantic-Fidelity Token Optimization
We first present the disadvantages of previous methods from the semantic-fidelity control. Then, we introduce our optimization strategy, including the motivation for feature disentanglement and the details of obtaining feature pairs. Finally, we present our training loss for optimization.